how do i solve this equation 1/m + w = c/m?
Answer by checkley77(12844)

1/m + w = c/m
1/M-C/M=-W
(1-C)/M=W
MW=1-C
M=(1-C)/W ANSWER.
W=(1-C)/M ANSWER.
-C=MW-1
C=1-MW ANSWER.